OpenStudy (anonymous):

A rectangular dog run is to be fenced so that one of the sides has to be 20 feet longer than an adjacent side(s). Which of the following can be used to represent the area of the dog run? A = s^2 A=s(s+20) A=s^2 + 20 A=2s + 2(s-20)

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Shalante

OpenStudy (jadedry):

one side = s. The other side = s+20 Therefore s * (s+20) = s^2 +20s = s(s+20)

OpenStudy (anonymous):

thank u